1. Please use the sets below to find the following (3pts each)
U = {a, b, c, d, e, f, g, h}
A = {a, b, c, d}
B = {c, e, g}
C = {d, e, f, g, h}

(a) A ∪ (B ∩ C)
(b) B ∪ C
(c) (B ∪ C)'
(d) A − (B ∪ C)
(e) (B ∪ C) − A
1
Expert's answer
2020-10-12T19:02:17-0400

We are given:

"U = \{ a, b, c, d, e, f, g, h \}"

"A = \{ a, b, c, d \}"

"B = \{ c, e, g \}"

"C = \{ d, e, f, g, h \}"

Now:

"B \cup C = \{ c,e,g \} \cup \{ d, e, f, g, h \}"

"\hspace{1 cm} \, = \{ c, d, e, f, g, h \}"

And:

"B \cap C = \{ c,e,g \} \cap \{ d, e, f, g, h \}"

"\hspace{1 cm} \, = \{ e, g \}"

a.)

"A \cup \left( B \cap C \right) = \{ a,b, c, d \} \cup \{ e, g \}"

"\hspace{1 cm}\hspace{0.9 cm} \, = \{ a, b, c, d, e, g \}"

b.)

"B \cup C = \{ c,e,g \} \cup \{ d, e, f, g, h \}"

"\hspace{1 cm} \, = \{ c, d, e, f, g, h \}"

c.)

"\left( B \cup C \right)^{'} = U - \left( B \cup C \right)"

"\hspace{1 cm} \hspace{0.4 cm} = \{ a, b, c, d, e, f, g, h \} - \{ c, d, e, f, g, h \}"

"\hspace{1 cm} \hspace{0.4 cm} = \{ a, b \}"

d.)

"A - \left( B \cup C \right) = \{ a, b, c, d \} - \{ c, d, e, f, g, h \}"

"\hspace{1 cm} \hspace{1 cm} = \{ a, b \}"

e.)

"\left( B \cup C \right) - A = \{ c, d, e, f, g, h \} - \{ a, b, c, d \}"

"\hspace{1 cm} \hspace{1 cm} = \{ e, f, g, h \}"